Cantona Manufacturing Co. purchased a machine on July 1, 2010, for $45,0 estimated five year life with a salvage value of $5,Could you solve clearly? Thanks a lot!!
Answer #1

Answer = E) $23,400

Double Declining Balance Rate = 1/Life *2 =(1/5)*2 =40%

Depreciation of 2010 = 45,000*40%*6/12 = 9,000

Depreciation of 2011= (45000-9,000)*40% = 14,400

Accumulated Depreciation = 9,000+14,400 = $23,400 (Answer)